OpenStudy (anonymous):

find f(x) when f(-2) = 10 for the anti-derivative (it like a squigly thing ) of (8x^3 -2x) dx

jimthompson5910 (jim_thompson5910):

So this you mean? \[\Large \int (8x^3-2x)dx\] this is also called the "integral"

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

the integral is really the opposite of the derivative for example if y = x^2, then dy/dx = 2x if I told you that dy/dx = 2x and you didn't know that y = x^2, then you'd just think backwards to go from dy/dx = 2x to y = x^2. BUT y = x^2 and any other function of the form y = x^2 + C will have the derivative of dy/dx = 2x. Examples: y = x^2 + 1, y = x^2 + 5, etc all have derivatives dy/dx = 2x so that's why integrals are also referred to as "antiderivatives". They undo derivatives.

jimthompson5910 (jim_thompson5910):

can you think of something that may derive to 8x^3 ? ie if I told you that dy/dx = 8x^3, then what could y be?

OpenStudy (anonymous):

2x^4 ?

jimthompson5910 (jim_thompson5910):

very good you can use a guess and check method, or you can use this rule here \[\Large \int (x^n)dx = \frac{x^{n+1}}{n+1}+C\] Notice how deriving the right side will get you to x^n

jimthompson5910 (jim_thompson5910):

also, a lot of derivative rules also apply to integrals as well for instance, with differentiation,we can pull out the coefficient and ignore it for the time being while we derive. The same is true with integration so what I'm saying is that we can do this \[\Large \int(8x^3)dx = 8*\int(x^3)dx\] and then use the rule above

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so far I am understanding

OpenStudy (triciaal):

great

jimthompson5910 (jim_thompson5910):

Using that rule and other integral rules, we get this \[\Large \int (8x^3-2x)dx = \int(8x^3) + \int(-2x)dx\] \[\Large \int (8x^3-2x)dx = 8\int(x^3) - 2\int(x)dx\] \[\Large \int (8x^3-2x)dx = 8\int(x^3) - 2\int(x^1)dx\] \[\Large \int (8x^3-2x)dx = 8(\frac{x^{3+1}}{3+1}+C) - 2(\frac{x^{1+1}}{1+1}+D)\] \[\Large \int (8x^3-2x)dx = 8*\frac{x^{4}}{4}+8C - 2*\frac{x^{2}}{2}-2D\] \[\Large \int (8x^3-2x)dx = 2x^4 - x^2 + C\] Note: the C and D constants combine to form one constant, and we usually just call it C at the very end

OpenStudy (anonymous):

yeah you always add that because you do not know if there is a constant or not

jimthompson5910 (jim_thompson5910):

To check the answer, you would derive 2x^4 - x^2 + C and you will get 8x^3 - 2x. So that confirms the antiderivative

jimthompson5910 (jim_thompson5910):

I'm assuming your book made F(x) the original function and f(x) the antiderivative ie, y = f(x), so dy/dx = F(x)

jimthompson5910 (jim_thompson5910):

that means \[\Large \int F(x)dx = f(x) + C\]

OpenStudy (anonymous):

but what do I do with the f(-2) = 10 part ? because that was the part that confused me

jimthompson5910 (jim_thompson5910):

so essentially we know that f(x) = 2x^4 - x^2 + C we also know that f(-2) = 10, so plug in x = -2 f(x) = 2x^4 - x^2 + C f(-2) = 2(-2)^4 - (-2)^2 + C then replace f(-2) with 10, since the two are equal, to get f(-2) = 2(-2)^4 - (-2)^2 + C 10 = 2(-2)^4 - (-2)^2 + C now solve for C and tell me what you get

OpenStudy (anonymous):

ok give me a moment

OpenStudy (anonymous):

-18

jimthompson5910 (jim_thompson5910):

so this particular f(x) function is really \[\Large f(x) = 2x^4 - x^2 - 18\]

OpenStudy (anonymous):

oh wow , that was so simple thank you so much for your help

jimthompson5910 (jim_thompson5910):

you're welcome